# AT-MoE: Adaptive Task-planning Mixture of Experts via LoRA Approach

### v1

@@ 3 AT-MoE: Methodology
In specific implementation, we need to maintain two type of trainable matrices. The first matrix is the group routing vector $W_{G}$, with a size of $N_{dim}*N_{G}$, where $N_{dim}$ is the input’s embedding size. Suppose for all the $N_{G}$ groups, the maximum number of sub-experts in each group is $N_{M}$. In specific training, we denote the adaptive group transformation as a function $F_{G}$(.). The detail for $F_{G}$(.) is described as follows: we first map according to the input’s embedding through $W_{G}$ to a grouped weight vector $W_{G}$ of size $1*N_{G}$, and then we perform temperature-based $SoftMax$ on $M_{G}$ the obtain the $W^{{}^{\prime}}_{G}$. The second matrix is the distribution routing matrix $W_{D}$, with a size of $N_{G}*N_{M}$. For those groups with less than $N_{M}$ experts, We directly pad the remaining weights with negative infinity, so that they do not participate in the $SoftMax$ calculation. We then apply the temperature-based $SoftMax$ operation on the $W_{D}$ transformed matrix for each group column and achieve the $W^{{}^{\prime}}_{D}$. Finally, we calculate the weighted sum from the dot product result between each task-specific LoRA matrix and the corresponding weight in $W^{{}^{\prime}}_{D}$. In addition, we use the merged training dataset for all the tasks to train an pre-merged LoRA representing a general-purpose expert. The output for all the experts is expressed as:

### v2

@@ 2 AT-MoE: Methodology
In specific implementation, we need to maintain two type of trainable matrices. The first matrix is the group routing vector $W_{G}$, with a size of $N_{dim}*N_{G}$, where $N_{dim}$ is the input’s embedding size. Suppose for all the $N_{G}$ groups, the maximum number of sub-experts in each group is $N_{M}$. In specific training, we denote the adaptive group transformation as a function $F_{G}$(.). The detail for $F_{G}$(.) is described as follows: we first map according to the input’s embedding through $W_{G}$ to a cross-group weight vector $W_{G}$ of size $1*N_{G}$, and then we perform temperature-based $SoftMax$ on $M_{G}$ the obtain the $W^{{}^{\prime}}_{G}$. The second matrix is the inside-group routing matrix $W_{D}$, with a size of $N_{G}*N_{M}$. For those groups with less than $N_{M}$ experts, We directly pad the remaining weights with negative infinity, so that they do not participate in the $SoftMax$ calculation. We then apply the temperature-based $SoftMax$ operation on the $W_{D}$ transformed matrix for each group column and achieve the $W^{{}^{\prime}}_{D}$. Finally, we calculate the weighted sum from the dot product result between each task-specific LoRA matrix and the corresponding weight in $W^{{}^{\prime}}_{D}$. In addition, we use the merged training dataset for all the tasks to train an pre-merged LoRA representing a general-purpose expert. The output for all the experts is expressed as: